OpenStudy (anonymous):

If the volume of a sphere is 18 cm cubed, what is the radius?

OpenStudy (anonymous):

V = 4/3 pi r^3. find r from that.

OpenStudy (anonymous):

Volume = 4/3 pi r^3 find r using that formula.

OpenStudy (anonymous):

omg but i need to know the radius.

OpenStudy (anonymous):

And you can find it, just work backwards. Plug in the volume you are given and solve for r.

OpenStudy (anonymous):

the equation for the volume of a sphere is (4*Pi*r)/3 solve for r by moving 3/(4*pi) to the other side of the equation so you get r^3 = (3*18)/(4*pi) simplify

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

if V = 4/3 pi r^3, then multiply both sides by 3/4 3/4V = pi r^3 divide both sides by pi r^3 = (3/4pi) V r = cube root of (3/4 pi) V, substitute V = 18 for your answer. Also, learn manipulation of equations. http://www.chem.tamu.edu/class/fyp/mathrev/mr-algeb.html

OpenStudy (anonymous):

khanacademy.org has good videos demonstrating how to do this.